XYZ Company enters into an exchange transaction with LMN Company. XYZ Company exchanges a building from one of its retail locations to LMN Company and ABC receives land to develop an amusement park. The building from ABC has a FMV of $385,000 and an original purchase price of $310,000 and depreciation taken of $65,000. The land from LMN has a basis and FMV of $295,000 and they give $90,000 cash to ABC.

  1. Does this exchange qualify for a like-kind exchange treatment?
  2. What is the realized gain on the exchange to ABC Company?
  3. What is the recognized gain on the exchange to ABC Company?
  4. What is the total basis in the property and cash received by ABC Company?
  5. What is the recognized gain to LMN Company on the transaction?
  6. What is the basis in the property received by LMN Company in the exchange?
Answer #1

A. In business when any asset is sold or exchanged the value considered is the actual value of the aaset.

In this case ABC exchange its building with LMN ,where

ABC COMPANY

the fair market value of  is - 385000

original price - 310000

less: depreciation- 65000

so the actual price at the time of exchange is - 245000

LMN COMPANY

the fair market value is -295000

add: cash paid- 90000

total value at the time of exchange- 385000

therefore the exchange does not qualify in like kind of exchange as LMN COMPANY pays full 385000 as fair price but receives an asset only of value 245000.

B. THE REALIZED GAIN ON EXCHANGE TO ABC COMPANY IS = 385000 - 310000 = 75000

realized gain is the profit that you gain with the difference between fair market value and the price that you sell your asset and it is not taxable.

C. THE RECOGNIZED GAIN ON EXCHANGE TO ABC COMPANY = 385000 - 245000 =140000

the recognized gain is the gain while selling your asset as is taxable. here ABC gives its asset for value of 245000 and received as asset valued 385000.so the asset received by ABC is of 140000 more that its asset given to LMN.

D. The Basis is 245000 and cash received is 90000.

E. There is no recognized gain to LMN because it receives less value towards its asset exchanged.

F. The basis received by LMN COMPANY IS 245000.

Here they have their asset valued at 295000 and ABC gives there asset valued at 245000 so 245000 is the basis.